# Water-based Quantum Dots Liquid Scintillator for Particle Physics

### v2

@@ 4.2 Simulation
The geometry created is a simplified replica of the physical setup taking into consideration the distance, size, and shape of the objects. Fig. 8 left shows an event display of our simulation in GEANT4. The PMT hemispherical glass surface is set as the sensitive detector. The WbQD material is treated as a liquid scintillator with a similar emission spectrum, located inside of a cylindrical glass. Utilising the necessary GEANT4 libraries such as electromagnetic processes, the physics processes can be constructed. Muons are generated randomly and they penetrate the WbQDs volume and produce scintillation photons that are detected by the PMT. The trajectories of muons are calculated by the solid angle defined by 2 cosmic ray taggers, and for simplicity all muons are 4 GeV in this simulation. In the simulation, we do not simulate PMT responses, instead, all photons which hit the surface of the PMT are registered with constant $15\%$ efficiency and the gain is $1.0\times 10^{7}$ at $-1100$ V according to our calibration.
Fig. 8 right shows data-MC comparison of their charge distributions above the pedestal. Here, the number of scintillation photons in the simulation is determined by adjusting the two distributions to match. A good agreement between experimental and simulated data is achieved around $4000$ photons/MeV, although with some deviations in the low charge region. This is likely due to the contamination of background events in the data, and further investigations are required to confirm this. The simulation adopts a simple geometry, and objects in the dark box, except those shown in Fig. 8, left, are ignored. However, we believe they do not affect the time and charge distributions much, as major signals are direct light travelling from the WbQD scintillator to the front of the PMT. Importantly, this is the first demonstration of the water-based quantum dots liquid scintillator with radiation detection; furthermore, the number of photons recorded here is promising for applications in particle and nuclear physics.
Figure 8: Simulation of the experimental setup in Geant4. (a) Event display of the simulation. Cosmic ray distributions are based on the configuration of 2 cosmic ray taggers and example muon trajectories are shown in green. (b) Data-simulation charge distribution comparison. Data pedestal is cut, and the number of scintillation photons per MeV is adjusted so that the distributions agree.
@@ 5 Discussions
The results of the simulations indicate that several thousands of photons are created by CdSe/ZnS QDs per 1 MeV of energy deposited, which is a surprisingly high number considering the low concentration of the dispersion ($42\pm 1$ nM). For reference, a typical organic fluor in a liquid scintillator, such as 2,5-diphenyloxazole (PPO), would be present at a concentration of $4-30$ mM and would be expected to produce around 10,000 photons from 1 MeV energy deposit. The larger physical cross-sections of the QDs ($\approx 20~{}\mathrm{nm^{2}}$ in our case), when compared to organic fluors ($0.5-1~{}\mathrm{nm^{2}}$), would be expected to somewhat mitigate for the lower concentration, but the differences in sizes are not sufficient to fully account for the observed photon rates. Furthermore, the PLQY of our QDs is also lower to that of a typical organic fluor ($9.5\%$ for WbQD sample and $85\%$ for PPO).
@@ 6 Conclusions
In this article, we investigate the potential of water-based quantum dots liquid scintillator for particle physics application. We first performed a phase transfer to CdS/ZnS QDs from hydrophobic phase to hydrophilic phase. Absorbance and emission peaks for CdS QDs are $\sim 470$ nm and $\sim 460$ nm for both in toluene and in water. These values fit in the range provided by the manufacturer, and we confirmed optical properties are mostly preserved in water-based quantum dots.
Photoluminescence quantum efficiency of the 2023 WbQD sample is measured to be $9.5\pm 0.5\%$. The value is lower than that provided by the manufacture, suggesting the loss due to phase transfer. Measured PLQY of other samples varies, and values go up or down over time and further study is needed to understand.
Hydrodynamic size of the toluene and water QD samples are measured to be $25\pm 9$ nm and $71.6\pm 0.3$ nm. TEM confirms the QD diameter to be $5.0\pm 1.6$ nm, and a consistent diameter is obtained from an empirical equation using the absorption spectrum. Large WbQD size is attributed to the oleic acid layer with polar water molecules. This layer takes over 1 year to reorganize itself and it is stable over 2 years. Furthermore, there is no sign of agglomeration. The phase transfer has a significant impact for both PLQY and QD size, and it is a subject to be optimized in future.
Finally, we performed scintillation measurements from the cosmic rays (atmospheric muons). The time distribution confirms the fast decay ($<5$ nm) of QD is preserved in water, and charge distribution shows relatively high scintillation yield. More quantitative studies are needed to confirm these numbers.
In conclusion, water-based quantum dots liquid scintillator is a promising new technology for particle physics. We have a special attention for a water solution of quantum dots to detect inverse beta decay processes from nuclear fission reactor neutrinos, where numerous programs are running all over the world [24]. Although liquid scintillator detectors [63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74] are popular base designs, other technologies such as solid scintillator [75, 76, 77, 78] are also explored on top of water based detectors [16, 17, 18, 19, 20, 21, 22, 79, 80]. Furthermore, reactor neutrino detection from neutrino-electron elastic scattering [81, 82] and neutrino-nucleus elastic scattering are established [83, 84].

### v3

@@ 4.2 Simulation
The geometry created is a simplified replica of the physical setup taking into consideration the distance, size, and shape of the objects. Fig. 9 left shows an event display of our simulation in GEANT4. The PMT hemispherical glass surface is set as the sensitive detector. The WbQD material is treated as a liquid scintillator with close emission spectrum, located inside of a cylindrical glass. Utilising the necessary GEANT4 libraries such as electromagnetic processes, the physics processes can be constructed. Muons are generated randomly and they penetrate the WbQDs volume and produce scintillation photons that are detected by the PMT. The trajectories of muons are calculated by the solid angle defined by 2 cosmic ray taggers, and for simplicity all muons are 4 GeV in this simulation. In the simulation, we do not simulate PMT responses, instead, all photons which hit the surface of the PMT are registered with constant $15\%$ efficiency and the gain is $1.0\times 10^{7}$ at $-1100$ V according to our calibration.
Fig. 9 right shows data-MC comparison of their charge distributions above the pedestal. Here, the three simulations are presented with the different numbers of scintillation photons, and the agreement with the data is assessed from the tails of these distributions to match. A good agreement between experimental and simulated data is achieved around $4000$ photons/MeV, although with some deviations for low charge region. This is likely due to the contamination of pedestal events in the data, and further investigations are required to confirm this. The simulation adopts a simple geometry, and objects in the dark box, except those shown in Fig. 9, left, are ignored. However, we believe they do not affect the time and charge distributions much, as major signals are direct light travelling from the WbQD scintillator to the front of the PMT. Importantly, this is the first demonstration of the water-based quantum dots liquid scintillator with radiation detection; furthermore, the number of photons recorded here is promising for applications in particle and nuclear physics.
Figure 9: Simulation of the experimental setup in Geant4. (a) Event display of the simulation. Cosmic ray distributions are based on the configuration of 2 cosmic ray taggers and example muon trajectories are shown in green. (b) Data-simulation charge distribution comparison. Data pedestal is cut, and simulation shows three predictions with the different number of scintillation photons per MeV, 3000 photons/MeV (blue dotted), 4000 photons/MeV (red dashed), and 5000 photons/MeV (green dash-dotted).
@@ 5 Discussions
The results of the simulations indicate that several thousands of photons are created by CdS/ZnS QDs per 1 MeV of energy deposited, which is a surprisingly high number considering the low concentration of the dispersion ($42\pm 1$ nM). For reference, a typical organic fluor in a liquid scintillator, such as 2,5-diphenyloxazole (PPO), would be present at a concentration of $4-30$ mM and would be expected to produce around 10,000 photons from 1 MeV energy deposit. The larger physical cross-sections of the QDs ($\approx 32~{}\mathrm{nm^{2}}$ in our case), when compared to organic fluors ($0.5-1~{}\mathrm{nm^{2}}$), would be expected to somewhat mitigate for the lower concentration, but the differences in sizes are not sufficient to fully account for the observed photon rates. Furthermore, the PLQY of our QDs is also lower to that of a typical organic fluor ($9.5\%$ for WbQD sample and $85\%$ for PPO).
@@ 6 Conclusions
In this article, we investigate the potential of water-based quantum dots liquid scintillator for particle physics application. We first performed a phase transfer to CdS/ZnS QDs from hydrophobic phase to hydrophillic phase. Absorbance and emission peaks for CdS QDs are $\sim 470$ nm and $\sim 460$ nm for both in toluene and in water. These values fit in the range provided by manufacturer, and we confirmed optical properties are mostly preserved in water-based quantum dots.
Photo luminescence quantum efficiency of the 2023 WbQD sample is measured to be $9.5\pm 0.5\%$. The value is lower than manufacture number, suggesting the loss due to phase transfer. Measured PLQY of other samples varies, and values go up or down over time and further study is needed to understand.
Hydrodynamic size of the toluene and water QD samples are measured to be $25\pm 9$ nm and $71.6\pm 0.3$ nm. TEM confirms the QD diameter to be $6.4\pm 1.2$ nm, and a consistent diameter is obtained from an empirical equation using the absorption spectrum. Large WbQD size is attributed to the oleic acid layer with polar water molecules. This layer takes over 1 year to reorganize itself and it is stable over 2 years. Furthermore, there is no sign of agglomeration. The phase transfer has a significant impact for both PLQY and QD size, and it is a subject to be optimized in future.
Finally, we performed scintillation measurements from the cosmic rays (atmospheric muons). The time distribution confirms the fast decay ($<6$ ns) of QD is preserved in water, and charge distribution shows relatively high scintillation yield. More quantitative studies are needed to confirm these numbers.
In conclusion, water-based qunatum dots liquid scintillator is a promising new technology for particle physics. We have a special attention for a water solution of quantum dots to detect inverse beta decay processes from nuclear fission reactor neutrinos, where numerous programs are running all over the world [24]. Although liquid scintillator detectors [65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76] are popular base design, other technologies such as solid scintillator [77, 78, 79, 80] are also explored on top of water based detectors [16, 17, 18, 19, 20, 81, 82]. Furthermore, reactor neutrino detection from neutrino-electron elastic scattering [83, 84] and neutrino-nucleus elastic scattering are established [85, 86].